Suppose data about the average number of hours people spend on social media websites at work is randomly collected for 5 days:

4, 4.5, 2.5, 2, 6

A.) If we assume this population is normally distributed, construct an 80% confidence interval for the population average amount of time people spend on social media websites at work.

B.) Suppose I want to know more about how workers spend their time throughout the day. I want to test the claim that people spend no more than 3 hours a day during their workday on Social Media. Test this claim (significance level =0.05).

(please show all work)
Answer #1

(A)

From the given data,the following statistics are calculated:

n = 5

\bar{x} = 19/5 = 3.8

s =1.6047

\alpha = 0.20

df = 5 - 1 = 4

From Table, critical values of tC = \pm 1.533

Confidence Interval:

\bar{x}\pm \frac{t_{C}\times s}{\sqrt{n}}=3.8\pm \frac{1.533\times 1.6047}{\sqrt{5}}=3.8\pm 1.1=(2.7,4.9)

So,

Answer is:

(2.7, 4.9)

(b)

H0: Null Hypothesis: \mu \leq 3 ( people spend no more than 3 hours a day during their workday on Social Media ) (Claim)

HA: Alternative Hypothesis:\mu > 3 ( people spend more than 3 hours a day during their workday on Social Media )

\bar{x} = 3.8

s = 1.6047

n = 5

\alpha = 0.05

df = n - 1 = 5 - 1 = 4

One Tail - Right Side Test

From Table, critical value of t = 2.132

Test Statistic is given by:

t=\frac{\bar{x}-\mu _{0}}{s/\sqrt{n}}=\frac{3.8-3}{1.6047/\sqrt{5}}=1.115

Since calculated value of t = 1.115 is less than critical value of t = 2.132, the difference is not significant. Fail to reject null hypothesis.

Conclusion:
The data support the claim that people spend no more than 3 hours a day during their workday on Social Media .